Mathematics 134 – Intensive Calculus for Science 2 Quiz 5 – March 3, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No calculators on this quiz.

A) (10) The LEFT(4), RIGHT(4), MID(4), and TRAP(4) methods are used to approximate ∫_{0}^{4} f(x) dx for the function below. Which method will give the largest numerical value? Which will give the second-largest?

Answers:

Largest:

Second-Largest:

B) (15) Find the volume of the solid of revolution obtained by rotating the region bounded by y = x^{2}, the x-axis, x = 0 and x = 2 about the line y = −1.